Based on a predicted level of production and sales of 29,000 units, a company anticipates total variable costs of $113,100, fixed costs of $29,000, and operating income of $203,580. Based on this information, the budgeted amount of operating income for 26,000 units would be: I know the correct answer is $179,520, but I don't know how they got that.

Multiple Choice

a. $179,520.

b. $208,520.

c. $66,420.

d. $101,400.

e. $309,920.
Answer #1
Total contribution magin at 29,000 units = Fixed costs+Operating income = 29000+203580= $232580
Contribution margin per unit=232580/29000= $8.02
Operating income for 26,000 units=(26000*8.02)-29000= $179520